OpenStudy (anonymous):

Quick math problem. 1/6+1/x=1/4 solve for x?

OpenStudy (anonymous):

what /6 = 1/4?

OpenStudy (anonymous):

\[\frac{ 1 }{ 6 }+\frac{ 1 }{ x }=\frac{ 1 }{ 4 }\] This is how it should look...

zepdrix (zepdrix):

Hey Kitty :) Fractions are no fun. I would recommend multiplying both sides of the equation by the LCM of each side. Hmmm, we have a 6, a 4, and an x. How bout we multiply both sides by 12x?

OpenStudy (anonymous):

So... how exactly would I write that out? I'm horrible at math.. sorry. :/

OpenStudy (anonymous):

boobies

zepdrix (zepdrix):

\[\Large\rm 12x\left(\frac{ 1 }{ 6 }+\frac{ 1 }{ x }\right)=\left(\frac{ 1 }{ 4 }\right)12x\]So here, we're multiplying each side by 12x. Every fraction is going to lose a little bit of something, but it will get rid of EVERY denominator, which is what we want.

zepdrix (zepdrix):

On the left side, we'll have to distribute the 12x to each term, \[\Large\rm 12x\cdot\frac{1}{6}\]What happens with the first term? Do you see how it will simplify?

OpenStudy (anonymous):

Will it be 2x ??

zepdrix (zepdrix):

Simplifying the 12 and the 6 in the bottom? Good good good.

zepdrix (zepdrix):

How bout the other two terms? :) Wanna give those a shot?

zepdrix (zepdrix):

\[\Large\rm 12x\left(\frac{ 1 }{ 6 }+\frac{ 1 }{ x }\right)=\left(\frac{ 1 }{ 4 }\right)12x\] \[\Large\rm 12x\cdot \frac{1}{6}+12x\cdot\frac{1}{x}=12x\cdot\frac{1}{4}\] \[\Large\rm 2x+12x\cdot\frac{1}{x}=12x\cdot\frac{1}{4}\]That's where we're at so far.

OpenStudy (anonymous):

Okay so \[12x * \frac{ 1 }{ x } = 12? \] and \[ 12x * \frac{ 1 }{ 4 } = 3x?\]

OpenStudy (anonymous):

oh

zepdrix (zepdrix):

\[\Large\rm 2x+12=3x\]Mmm ok great job! :) Much easier to solve now, yes?

OpenStudy (anonymous):

Yes!! So i just divide 2x from both sides now?

OpenStudy (anonymous):

So I can get x on one side..?

zepdrix (zepdrix):

Hmm 2x is being ADDED to the left side. Do the inverse operation. Inverse of addition is not division! :D

OpenStudy (anonymous):

Aha... I subtract 2x from both sides...

zepdrix (zepdrix):

cool c:

OpenStudy (anonymous):

So then x is equal to 12!

OpenStudy (doc.brown):

\[\frac{1}{6}+\frac{1}{x}=\frac{1}{4}\]\[\frac{1}{6}-\frac{1}{6}+\frac{1}{x}=\frac{1}{4}-\frac{1}{6}\]\[\frac{1}{x}=\frac{1}{4}-\frac{1}{6}\]\[\frac{1}{x}=\frac{6}{6}\cdot\frac{1}{4}-\frac{4}{4}\cdot\frac{1}{6}\]\[\frac{1}{x}=\frac{6}{24}-\frac{4}{24}\]

zepdrix (zepdrix):

Yay good job \c:/ Meow meow kitty meow!

OpenStudy (anonymous):

You rock! Thank you so much!!

OpenStudy (doc.brown):

\[\frac{1}{x}=\frac{6}{24}-\frac{4}{24}\]\[\frac{1}{x}=\frac{2}{24}\]\[\frac{1}{x}=\frac{1}{12}\]\[\frac{1x}{x}=\frac{1x}{12}\]\[1=\frac{x}{12}\]\[12=\frac{12x}{12}\]\[12=x\]

OpenStudy (anonymous):

now suck him off